OpenStudy (henrietepurina):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (henrietepurina):

y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

hartnn (hartnn):

can you find the slope of 3x - y = -3 ?

OpenStudy (henrietepurina):

hmm... im not very good at that

hartnn (hartnn):

try to isolate 'y' 3x-y =-3 so, -y = -3x-3 or y = 3x+3 got this ?

hartnn (hartnn):

then compare it with y=mx+c where m is the slope and you'll get the slope of this line